# Towards Internet-Scale Training For Agents

### v1

@@ 6.2 Improving Generalization
Figure 11: Our data improves generalization. We train agents with all human data from the WebLINX and Mind2Web training sets, and resulting agents struggle to generalize to more diverse test data. Adding our data improves generalization by +149.0% for WebLINX, and +156.3% for Mind2Web.
To understand how agents trained on data from our pipeline generalize to diverse real-world sites, we construct a more diverse test set than Mind2Web and WebLINX using 500 held-out demonstrations produced by our pipeline. Shown in Figure 11, we train agents using all human data in the training sets for WebLINX and Mind2Web, and compare the performance with agents trained on 80% human data, and 20% data from our pipeline. Agents trained with our data achieve comparable performance to agents trained purely on human data on the official test sets for the WebLINX and Mind2Web benchmarks, suggesting that when enough human data are available, synthetic data may not be necessary. However, when evaluated in a more diverse test set that includes 500 sites not considered by existing benchmarks, agents trained purely on existing human data struggle to generalize. Training with our data improves generalization to these sites by +149.0% for WebLINX agents, and +156.3% for Mind2Web agents, with the largest gains in generalization Step Accuracy appearing for harder tasks.
@@ 7 Conclusion
We present a pipeline that unlocks new scaling potential for language model agents by generating internet-scale web navigation data for diverse sites. We employ pretrained language models to generate, attempt, and evaluate web navigation tasks at scales beyond what humans can manually annotate. Our results show that language models outperform human annotators at detecting safe websites, and proposing feasible web navigation tasks. In addition, we find that language models can judge successful web navigation trajectories with an accuracy up to 93.1% for their most confident predictions. We scale task generation to the top 1M sites in the CommonCrawl PageRank, and test agents on 150k live web navigation tasks. Agents based on Llama 3.1 70B solve 16.7% of these tasks zero-shot. In data-limited settings derived from Mind2Web and WebLINX, we improve Step Accuracy by up to +89.5% and +122.1% respectively for agents trained on mixtures of data from our pipeline, and human data. When training agents with all available human data from these benchmarks, agents fail to generalize to diverse real sites, and adding our data improves their generalization by +149.0% for WebLINX and +156.3% for Mind2Web.
Our work reveals several exciting directions for future work. First, our work can be scaled further. The latest CommonCrawl release contains data for more than 300 million sites, suggesting another 1,000 times more data than we explored could be available for training agents by scaling the pipeline. In addition, our language model judge was employed offline, and its high accuracy suggests that it could be used to guide an online algorithm. Finally, we considered primarily text-based agents in this work, and our pipeline could be extended to generate data for multimodal tasks.
@@ 8 Acknowledgments
The authors recognize funding via an unrestricted gift from Amazon. In addition, the leading author was partially funded by the US Department of Defense via the NDSEG Fellowship. We thank fellow Ph.D. students who provided feedback on versions of this manuscript, including Michael Luo, Justin Wong, Tiffany Min, Murtaza Dalal, Jing-Yu Koh, Boyuan Chen, and others. We also thank faculty who provided feedback and direction for parts of this work, including Aviral Kumar, Graham Neubig, Shuyan Zhou, and others.
@@ Appendix A Limitations & Safeguards
Language model agents present unique challenges and risks when applied to live tasks on the internet. For instance, agents visiting shopping sites can influence the statistics produced by analytics tools, which can impact prices on products, and product decisions from companies. Furthermore, agents visiting harmful content can add such harmful content to datasets, and perpetuate harmful behaviors into the training data for future agents. We mitigate these risks by carefully designing the task proposal stage of the InSTA pipeline. We consider the risks posed to analytics tools by limiting the engagement between agents and sites. We generate only one task per website, and we limit agents to just 10 actions per site, which includes clicks, typing, dropdown selection actions, and more. By limiting the interaction between agents and sites, the change in website traffic generated by the InSTA pipeline is minimal (just $30$ seconds of interaction per site on average). By utilizing data from the InSTA pipeline in an offline fashion, as in Section 6 of the main paper, no additional web traffic is generated when training agents. To ensure that agents do not modify the state of the web (i.e. avoid attempting to make purchases, avoid leaving comments on posts, avoid making accounts, etc), we provide an instruction in the system prompt of the task proposer (see Figure 2) to avoid writing tasks that require the agent to modify the state of the web.
The task proposer is instructed via the system prompt to filter out sites with harmful content, sites not intended for user access, and sites that require making an account to operate (such as social media, and forums). We explore the performance of the task proposer at filtering out unsuitable sites in Section 4.2, and find that all models detect unsuitable sites with a recall from $0.98$ to $1.0$, and accuracy up to 97%, suggesting our filter is reliable. Sites used to benchmark the performance of the safety filter are discussed in Appendix E, and include harmful, and mature content.
@@ Appendix B Ethical Considerations
One important ethical consideration when gathering data from the internet is to handle copyrighted, private, and sensitive materials carefully. The internet contains vast amounts of personal data created by users that includes personally-identifying-information that should not be included in public datasets. We address this ethical consideration in two ways. First, the task proposer is instructed to filter out social media sites and forums that are likely to contains personally-identifying-information. Second, we store and release only the prompts we used, and traces for agents’ actions—importantly, we do not release any web source code that could be used to recover sensitive data. These steps significantly reduce, but do not completely eliminate the risk that private, and sensitive materials are included in our data, and methods for detecting, replacing, and removing such materials from datasets remains an important task for researchers working on safety.
@@ Appendix C Broader Impacts
As their capabilities broaden, language model agents are being increasingly used to operate real-world systems and APIs. This shift comes with several benefits and risks. Agents that operate your computer to aid in work tasks can significantly boost productivity for certain workers, but can displace others whose jobs have been fully automated. Agents that operate web browsers to complete personal tasks for users can provide convenience, but expose a new attack vector where compromised agents perform unintended actions. Certain risks can be mitigated with proper safeguards, such as post-processing data to prevent jail-breaking, but other risks are existential in nature, and harder to address purely from a research perspective.
Our data pipeline aims to facilitate internet-scale training for agents using offline data. Training from offline data is important because current agent capabilities are low, and training them on live websites risks taking bandwidth away from real users, while polluting the internet with noisy LLM outputs. Due to these concerns, we advocate for researchers building internet-scale agents to follow the guidelines set forth in this work: (1) agents should not modify the state of the internet by making accounts, creating posts, or submitting content; (2) agents should be throttled to a maximum rate of interaction, and a maximum number of interactions per website; (3) copyrighted, private, and sensitive materials should be removed from training data.
@@
Akin to robots.txt directives, website creators should have a standard format to specify how internet agents are allowed to interact with their websites—if at all. Desireable controls include rate limits for interactions, limits for maximum numbers of interactions, restrictions to allow agents to interact with certain pages and not others, and restrictions on the kind of data on webpages that agents are allowed to observe (achieved via tagging elements to hide their content from agents). In addition to restricting the data available to agents, website creators should have the ability to specify locations for “playgrounds” that replicate certain key functions of their site with virtual tasks and simulated data that are intended to teach agents how to operate their site while directing traffic from agents away from their primary user-facing platform.
@@ Appendix E More Details On Task Generation
Figure 12: System prompt for task generation. We carefully design the system prompt for task generation to ensure that sites not suitable for inclusion in the training data for agents are detected and removed. This prompt ensures that proposed tasks are passive in nature, and only involve retrieving information—active tasks like making posts and creating accounts are explicitly not allowed.
We provide the system prompt used for task generation in Figure 12. This prompt was provided to Llama 3.1 70B, GPT-4o, and Gemini 1.5 Pro to generate tasks and filter sites unsuitable for annotation in Section 4. We carefully designed this system prompt to enforce that generated tasks are passive in nature, and do not modify the state of content on the internet. In addition to this system prompt, we employed a list of 100 hand-picked in-context examples of website URLs and appropriate tasks, which are provided in the following JSON list. When querying an LLM to generate tasks, we randomly sample 16 in-context examples from the following list, and provide only these examples to the LLM. This helps promote diverse tasks.
@@ E.1 Details For Safety Experiment
This list of examples is also provided in our code release, alongside the script that we used to generate task proposals for the top 1M sites in the CommonCrawl PageRank (The Common Crawl Foundation, 2024). Using these prompts for task generation, we can filter our sites that are unsuitable for annotation, due to containing harmful content, or sensitive user data. To evaluate the performance of our filter, we employed a set of 100 curated websites, where 50 are manually verified as safe, and 50 are manually verified as unsafe based on the filtering conditions. These sites were chosen to span popular sites that typical annotators are likely familiar with, and less popular sites that annotators may not already be familiar with.
@@
We provide the list of safe sites in the preceding code block, and in the code release, but we provide the list of unsafe sites in just the code release to keep this manuscript clean of harmful content. Using these lists, we conduct an experiment measuring the accuracy, precision, and recall of our safety filter for detecting sites that are not suitable for training agents. In this experiment, we run the task proposer with the same prompts used to scale annotation the top 1M sites in the CommonCrawl PageRank, and we consider a site to be marked positive for unsafe content if the task proposer generates “N/A” rather than predicting a task. To calculate the human baseline for this experiment, we showed participants the system prompt observed by the task proposer, and asked participants to label sites as safe or unsafe using this per-example prompt:
[…]
Human participants were not allowed to visit the URL shown, and had to determine whether the site is safe for annotation purely from their prior knowledge (the same conditions faced by the task proposer). The 100 sites for the safety experiment were shuffled into a uniformly random order to ensure the order of annotation did not bias the human annotators predictions. One human participant was used to obtain the human baseline result in Table 3.
@@ E.2 Details For Reliability Experiments
Similar to the previous safety experiment, we employed human participants to obtain a human baseline for task feasibility. In particular, we showed human participants the system prompt in Figure 12 for the task proposer, and had them write a task for each of the following websites without visiting the URL (the same conditions faced by the task proposer). The following 100 sites were shuffled into a uniformly random order to ensure the participants were not influenced by the order in which sites were shown. After tasks were proposed by participants, and by LLMs, we evaluated the expert feasibility of tasks by manually attempting to complete the tasks proposed by each set of participants, and marking tasks as feasible, or not feasible based on our own ability to complete them. In total, we annotated 400 tasks, which required 8 hours of annotation. One human participant was used to obtain the human baseline result in Table 4.
@@ E.3 Automatic Task Categorization
To better understand the statistics of generated tasks, we employ Llama 3.1 70B to assign task categories. We prompt Llama 3.1 70B with the system prompt in Figure 13 to assign a category in 3 words or less to encourage simple categories. Categories have 16.9 tasks on average, and 953 categories have more than the mean, while 7741 have less than the mean. There is occasional overlap between categories, which can be observed in Figure 14, but for the purposes of understanding performance by category, overlap is acceptable provided categories have sufficiently large numbers of tasks, and performance per category can be accurately calculated. We provide our task categorization script in the official code release.
@@ Appendix F Understanding Agent Capabilities & Limitations
Figure 14: Largest categories for internet-scale task generation. We assign categories to 150k web navigation tasks generated by our pipeline in Section 4, and visualize the number of tasks for each of the largest 70 categories. Top categories include article search, news search, recipe search, product lookup, and more. The top 12 task categories have more than 1600 tasks assigned to each of them, the mean number of tasks per category is 16.9, and 89% of categories (7741 in total) have fewer than the mean number of tasks.
Figure 15: Most successful categories for internet-scale task generation. We explore the rates of task completion for the top categories of tasks generated by our pipeline. We restrict our focus to categories where at least 100 tasks are assigned, and plots the success rates for the top 70 of such categories. Results show that 22 categories are solved with more than a 50% rate with agents based on Llama 3.1 70B.
To complete the analyses presented in Section 5, we explore the categories of tasks that agents succeed at most frequently. Shown in Figure 15, we plot the average judge success probability prediction $\mathbf{r}_{T}$ versus task category for the top 70 most successful categories that have at least 100 tasks assigned to them. Based on the figure, top categories include search for contact information, finding hours of operation, looking up biographical information, obtaining current weather forecasts, and conducting health research. Based on these results, the top 22 categories are solved with more than a 50% rate using agents based on Llama 3.1 70B running zero-shot. As stronger models are developed, the success rates for agents running in our pipeline are likely to improve, and the quality of the data we generate will jointly improve.
In addition to studying the best-performing categories, we also explore the limitations of current agents via their least successful categories. Shown in Figure 16, we select the bottom 70 categories in terms of their average judged success probability for categories with at least 100 tasks assigned to them. Many of these categories require agents to remember and reason about previous interactions, such as the product comparison category. For this category, an agent must review several products, and compare their details from memory. In these cases, access to a note-taking app may improve performance. Additionally, certain task categories involve requests that are not feasible given the limitations of the Playwright API, including categories for downloading reports / manuals, and opening and playing files. While these tasks are not currently feasible, providing agents with a fully-operable virtual computer environment could unlock these abilities in future work.

### v2

@@ 6.2 Agents Transfer To New Domains
Figure 9: Our data transfers to static benchmarks. We train agents with all human data from the WebLINX and Mind2Web training sets, and resulting agents struggle to generalize to more diverse test data. Adding our data improves generalization by +149.0% for WebLINX, and +156.3% for Mind2Web.
Static Benchmarks.
To complement the online evaluation in Figure 8, we also explore how our data impacts agents trained on static benchmarks. We first train baseline agents on human demonstrations from the Mind2Web (Deng et al., 2023) and WebLINX (Lù et al., 2024) datasets. We write a preprocessor to convert our data into the expected action-observation formats these benchmarks use (which involves discarding our reasoning trace). With data converted, we compare the baseline to agents trained on a mix of 80% human data, and 20% our data, and test on (1) their official test set, and (2) 500 diverse sites from our official test set. Results in Figure 9 show that agents trained with our data perform equally well on the original test sets for these benchmark, but generalize better to our harder test set. Overall, we see +149.0% for WebLINX agents, +156.3% for Mind2Web agents, and gains in Step Accuracy on our test set are larger for the harder tasks. Additional experimental details are listed in Appendix H. On three popular benchmarks (WebVoyager, WebLINX, Mind2Web), our pipeline leads to capable agents and does not depend on human annotations.
6.3 Performance Scales With Test-time Compute
To understand the ability for web agents to scale with additional test-time compute, we ablate the number of tokens in the reasoning budget. Figure 10 shows the success rate as a function of the reasoning budget for Gemini 2.5 Flash, the top-performing agent we tested. There is a monotonic improvement in the success rate as the reasoning budget increases, and the trend suggests that performance may not be saturated with 500 reasoning tokens. Training language model agents to reason before taking actions is a promising path to better agents, and we are releasing a large reasoning dataset for multimodal agents to study this. Our dataset contains 2.2M screenshots, 2.2M reasoning traces for actions, 150k traces for judge evaluations, and led to the results in Section 6.1. The data will be linked on our website, alongside an official huggingface dataset for tasks.
Figure 10: Agents improve with a larger reasoning budget. We ablate the number of tokens in the reasoning budget for the top-performing agent, and see a monotonic improvement in the success rate as the reasoning budget increases. Gemini 2.5 Flash has a 70% success rate with a budget of 500 reasoning tokens, up from 60% for a budget of 50 tokens. The scaling of performance with the reasoning budget highlights a promising behavior in successful web agents that can be studied.
@@ 7 Conclusion
In the spirit of deep learning, we have developed an approach to efficiently harness internet data for LLM agents, and have unlocked a dynamic internet-scale environment. In building this environment, we presented a method to annotate 150k diverse sites with challenging agentic tasks, and showed how training on data from our pipeline allows small models to compete with frontier LLMs as agents, on a fraction of the budget. Our pipeline consists of a task proposer, agent, and judge driven by pretrained language models that together curate high-quality data for agents, without human intervention.
Our top checkpoint for Qwen 3 1.7B has a success rate of 56.9% on our test environment, outperforming the data collection policy Qwen 3 235B, beating the 235 times larger Llama 4 Maverick, and reaching 94.7% of the performance of Gemini 2.5 Flash, while being smaller and faster than these. Our models zero-shot transfer to WebVoyager, and scale with test-time compute. We are releasing the entire pipeline, including code, models and data, so that it may serve as a foundation for researchers to build the next generation of language model agents with internet data.
7.1 Future Work
Our work reveals several exciting directions in future work. First, our work can be scaled further: the latest Common Crawl release contains data for more than 300M sites, suggesting another 1,000 times more data could be available for agents by scaling the pipeline. In addition, we trained agents to optimize the judge scores indirectly via filtered SFT, and the judge’s high accuracy suggests that it could be optimized via reinforcement learning instead. RL is especially promising for how it can improve reasoning capabilities in agents. Finally, while the data we collect is multimodal, we focus on textual tasks in this paper, and our pipeline could be extended to produce multimodal tasks.
@@ 8 Acknowledgments
The authors recognize funding via an unrestricted gift from Amazon. In addition, the leading author was partially funded by the US Department of Defense via the NDSEG Fellowship. We thank fellow Ph.D. students who provided feedback on versions of this manuscript, including Michael Luo, Justin Wong, Tiffany Min, Murtaza Dalal, Jing-Yu Koh, Boyuan Chen, and others. We also thank faculty who provided feedback and direction for parts of this work, including Aviral Kumar, Graham Neubig, Shuyan Zhou, and others. Finally, we thank the open source community for developing systems and methods for serving LLMs efficiently, without which this paper would not be possible.
@@ Appendix A Limitations & Safeguards
Language model agents present unique challenges and risks when applied to live tasks on the internet. For instance, agents visiting shopping sites can influence the statistics produced by analytics tools, which can impact prices on products, and product decisions from companies. Furthermore, agents seeing harmful content on the web can add that content to datasets inadvertently, and propagate harmful behaviors to future agents. We mitigate these risks in the design of the task proposal stage. We consider the risks posed to analytics tools by limiting the engagement between agents and sites. We generate only one task per website, and we limit agents to just 30 actions per site, which includes clicks, typing, dropdown selection, and more. By limiting the interaction between agents and sites, the change in website traffic generated by the InSTA pipeline is minimal (just $90$ seconds of interaction per site on average). By utilizing data from the InSTA pipeline in an offline fashion, as in Section 6 of the main paper, no additional web traffic is generated when training agents. To ensure that agents do not modify the state of the web (i.e. avoid attempting to make purchases, avoid leaving comments on posts, avoid making accounts, etc), we provide instruct the task proposer (see Figure 2) to avoid writing tasks that require the agent to interact with personal data, or user accounts.
The task proposer is instructed via the system prompt to filter out sites with harmful content, sites not intended for user access, and sites that require making an account to operate, including social media, and forum sites. There is likely a manner to safely train agents to operate user accounts, but we leave this task to future researchers. We explore the performance of the task proposer at filtering out unsuitable sites in Section 4.2, and find that all models detect unsuitable sites with a recall from $0.98$ to $1.0$, and accuracy up to 97%, suggesting our filter is reliable. Sites used to benchmark the performance of the safety filter are discussed in Appendix E, and thoroughly test the safety filter.
To remove Personally Identifiable Information (PII) from the data used for training agents, we include scrubadub, an industry standard PII removal tool for python developed by Leap Beyond, a data consultancy based in the European Union. Our pipeline has an argument that toggles the usage of scrubadub to remove PII from all website data, and we recommend this option be set.
@@ Appendix B Ethical Considerations
One important ethical consideration when harnessing internet data is to carefully handle copyrighted, private, and sensitive materials. The internet contains vast amounts of PII, which should be avoided when training models. We address this in two ways. First, we instruct the task proposer to filter out sites that may contain PII, including social media websites, and forums. Second, our pipeline has an argument that toggles the usage of scrubadub to remove PII, and we recommend this option be set. To prevent copyrighted materials from being used for training, we do not allow the agent to copy any books, documents, videos, audio, or files. The information present in the HTML DOM is publically available, and is converted into a sparse Markdown summary that focuses on interactive elements, and significantly transforms the underlying webpage in a manner that is not reversible.
These steps significantly reduce, but do not completely eliminate the risk that private, and sensitive materials are shown to the agent, and methods for detecting, replacing, and removing such materials from text and images remains an important task for researchers working on safety.
@@ Appendix C Broader Impacts
As their capabilities broaden, language models are being used to interface with real-world systems. This shift comes with several benefits and risks. Agents that operate your computer to aid in work tasks can significantly boost productivity for certain workers, but can displace others whose jobs have been fully automated. Agents that operate browsers to complete personal tasks provide convenience, but expose a new attack vector where rogue agents perform unintended actions. Certain risks can be mitigated with proper safeguards, such as post-processing data to prevent jail-breaking, but other risks are existential, harder to address purely with new agent research, and may require policy changes.
@@
Akin to robots.txt directives, website creators should have a standard format to specify how internet agents are allowed to interact with their websites, and what information on webpages agents are allowed to see. Desireable controls include rate limits for interactions, limits for maximum numbers of interactions, restrictions to allow agents to interact with certain pages and not others, and restrictions on the kind of data on webpages that agents are allowed to observe (achieved via tagging elements to hide their content from agents). In addition to restricting the data available to agents, website creators should have the ability to specify locations for “playgrounds” that replicate certain key functions of their site with virtual tasks and simulated data that are intended to teach agents how to operate their site while directing traffic from agents away from user-facing pages.
@@ Appendix E More Details On Task Generation
Figure 11: System prompt for the exploration phase of task generation. We design the system prompt for task generation to detect and remove unsafe websites. This prompt ensures that tasks are passive, and do not modify content on a website. Refer to the next figures for the in-context examples used for the task proposer, and the system prompt used in the feedback step.
We provide the system prompt used in the first phase of the task generation loop in Figure 11. This prompt was provided to Llama 3.1 70B, GPT-4o, and Gemini 1.5 Pro to generate tasks and filter sites unsuitable for annotation in Section 4. We carefully designed this system prompt to enforce that generated tasks are passive, and do not modify content on a website. In addition to this system prompt, we employed a list of 100 hand-picked in-context examples of website URLs and appropriate tasks, which are provided in the following JSON list. When querying an LLM, we randomly sample 16 in-context examples from the list, and provide only these examples to the LLM to generate a task to guide exploration of the site. This improves diversity in the exploration phase.
@@
We also provide the system prompt used in the second phase of the task generation loop, where trajectories from agents are fed back to the task proposer, which generates a harder, grounded task. This prompt instructs the task proposer to create a challenging task based on how an expert user could be expected to use the shown website. The task proposer also predicts a list on intermediate steps that can be used as a hint for agents, and a success criteria that can be used to improve the verifier.
You are a helpful assistant designing tasks for a web automation script. I will show you previous runs of the script, including previous tasks, webpages, actions, and performance reviews, formatted in markdown. Help me design *challenging* new tasks.
## Formatting The Proposed Task
Format your task in the following JSON schema:
‘‘‘json
"proposed_task": str,
"steps": List[str],
"criteria": str
Here is what each key means:
- proposed_task‘: A specific, challenging task that an expert user might leverage this website to complete.
- Must not require making an account, logging in, submitting personal information, making a purchase, or placing an order.
- steps‘: Steps an expert user would follow to complete the proposed task.
- criteria‘: The required answer, and criteria to determine if the task was completed.
## Example Tasks For Inspiration
Suppose you want to design a task around the C-to-C Hose-Shut-Off Valve on awg-fittings.com’:
‘‘‘json
"proposed_task": "What is the C-to-C Hose-Shut-Off Valve length in mm?",
"steps": [
"Navigate to awg-fittings.com’",
"Open the product catelog for fittings",
"Locate the product listing for the C-to-C Hose-Shut-Off Valve",
"Find the product length in mm, and respond with that length in the answer"
"criteria": "The answer should include the specific length of ’237 mm for this product"
Suppose you want to design a task around the document The Angora cat; how to breed train and keep it on biodiversitylibrary.org’:
‘‘‘json
"proposed_task": "Open a scanned copy of The Angora cat; how to breed train and keep it’.",
"steps": [
"Navigate to biodiversitylibrary.org’",
"Search for The Angora cat; how to breed train and keep it in the search bar",
"Click on the title of the document in the search results",
"Confirm the correct document is displayed in an embedded PDF reader"
"criteria": "The final webpage should display the correct document in an embedded PDF reader"
Suppose you want to design a task around the Generative Adversarial Networks paper on scholar.google.com’:
‘‘‘json
"proposed_task": "How many citations does the paper Generative Adversarial Networks have?",
"steps": [
"Navigate to scholar.google.com’",
"Search for Generative Adversarial Networks in the search bar",
"Locate the correct paper in the search results",
"Find an up-to-date citation count, and respond with that count in the answer"
"criteria": "The answer should include an up-to-date citation count, which is ’80613’ as of April 2025"
Suppose you want to design a task around the word serendipity on wiktionary.org’:
‘‘‘json
"proposed_task": "What is the definition and etymology of the word serendipity’?",
"steps": [
"Navigate to wiktionary.org’",
"Search for serendipity in the search bar",
"Find the definition and etymology sections of the serendipity page",
"Summarize the contents of these sections in the answer"
"criteria": "The answer should mention Serendip (or Serendib), coined by English writer and politician Horace Walpole in 1754"
Thanks for helping me design challenging new tasks, please follow the instructions carefully. Start your response with an analysis for how an expert user would leverage this website, followed by a step-by-step breakdown of your proposed task, and finally, enter your task in the JSON format. Respond in 500 words.
@@ E.1 Details For Safety Experiment
Using these prompts for task generation, we remove unsafe websites. To evaluate the performance of our filter, we employed a set of 100 curated websites, where 50 are manually verified as safe, and 50 are manually verified as unsafe based on the filtering conditions. These sites were chosen to span popular sites that typical users are likely familiar with, and less popular websites that may be underrepresented in LLM training data.
@@
We provide the list of unsafe sites in the code to keep this manuscript clean. Using these lists, we conduct an experiment measuring the accuracy, precision, and recall of our safety filter for detecting sites that are not suitable for training agents. In this experiment, we run the initial exploration phase of the task proposer with the same prompts used in the main experiments (shown in the last section), and we consider a website to be marked positive for unsafe content if the task proposer generates “N/A” rather than a task. We then calculate accuracy, precision, and recall for various LLMs.
@@ E.2 Details For Reliability Experiments
We evaluated the verifiable rate of tasks generated by the initial phase of task generation manually. For a set of 100 randomly sampled websites marked as safe by the task proposer in the initial phase, we attempted to complete the generated task, and checked that enough information is present on the website that a solution can be verified. In total, we annotated 300 tasks for Table 2 in 6 hours, and provide the 100 website URLs used in this experiment in the following code block.
@@ E.3 Automatic Task Categorization
We employ Llama 3.1 70B to categorize tasks. We prompt Llama 3.1 70B with the system prompt in Figure 12 to assign a category in 3 words or less to encourage simple categories. Categories have 16.9 tasks on average, and 953 categories have more than the mean, while 7741 have less than the mean. There is occasional overlap between categories, which can be observed in Figure 13, but for the purposes of understanding performance by category, overlap is acceptable provided categories have sufficiently large numbers of tasks, and performance per category can be accurately calculated. We provide our task categorization script in the official code release.
@@ Appendix F Understanding Agent Capabilities & Limitations
Figure 13: Largest categories for task generation. We categorize 150k tasks generated by our pipeline in Section 4, and visualize the number of tasks in the largest 70 categories. Top categories include article search, news search, recipe search, and product lookup. The top 12 task categories have more than 1600 tasks assigned to each of them, the mean number of tasks per category is 16.9, and 89% of categories (7741 in total) have fewer than the mean number of tasks.
Figure 14: Most solved categories for task generation. We explore the completion rates for the top categories of tasks generated by our pipeline. We restrict our focus to categories where at least 100 tasks are assigned, and plots the success rates for the top 70 categories. Results show that 22 of these categories are solved with more than a 50% rate with zero-shot agents based on Llama 3.1 70B.
To complement the analyses presented in Section 5, we explore the categories of tasks that agents succeed at most frequently. Shown in Figure 14, we plot the average judge success probability prediction $\mathbf{r}_{T}$ versus task category for the top 70 most successful categories that have at least 100 tasks assigned to them. Based on the figure, top categories include searching for contact information, finding hours of operation, looking up biographical information, obtaining current weather forecasts, and conducting health research. Based on these results, the top 22 categories are solved with an average success probability $>0.5$ using zero-shot agents based on Llama 3.1 70B. As stronger models are developed, the success rates for agents running in our pipeline are likely to improve, and the quality of the data we generate will jointly improve.
In addition to studying the best-performing categories, we also explore the limitations of current agents via their least successful categories. Shown in Figure 15, we select the bottom 70 categories via their average judge success probability for categories with at least 100 tasks assigned. Many of these categories require agents to remember and reason about previous interactions, such as the product comparison category. For this category, an agent must review several products, and compare details from memory. In these cases, access to a note-taking tool may improve performance. Additionally, certain task categories involve requests that are not feasible given the limitations of the Playwright API, including categories for downloading reports / manuals, and opening files. While these tasks are not currently feasible, providing agents with a fully-operable virtual computer environment with applications pre-installed could unlock these abilities in future work.